If the 5-year discount factor is 0.91, the interest rate must be equal to ________
Answer #1

1 / (1 + R)5 = 0.91

(1 + R)5 = 1.0989

Apply 'Log' on both sides

Log (1 + R)5 = Log (1.0989)

5 * Log (1 + R) = 0.04095817

Log (1 + R) = 0.008191635

Apply 'Antilog' on both sides

Antilog (log (1 + R) = Antilog (0.008191635)

1 + R = 1.019041

R = 0.019041

R = 1.9041%